### v1

@@ 2.2.5 LCOGT
The Las Cumbres Observatory Global Telescope Network (LCOGT; Brown et al. 2013) observed GRB 240825A using the 1 m telescope. Observations of the afterglow of GRB 240825A began at 01:41:29 UT on August 26, 2024, 9.81 hours after the GBM trigger. A total of 10 observing runs were conducted between August 26 and 31, resulting in a set of B, V, R, and I band images. The exposure time of each image was set to 300 seconds, based on the trend of brightness estimated from other observations’ measurements and the limiting magnitude of the LCOGT telescopes. The photometry calibration is carried out using the USNO-B1 catalogues (Monet et al. 2003). The source is clearly detected by stacking the images from the first nine observations. All photometric data are tabulated in Table 7.
@@ 2.2.6 NOT
We triggered the 2.56-meter Nordic Optical Telescope (NOT) at the Roque de los Muchachos Observatory, La Palma, Spain, several times, utilizing the Alhambra Faint Object Spectrograph and Camera (ALFOSC). Photometric calibration was performed using the Pan-STARRS DR2 catalog (Flewelling 2018a). Photometric data are tabulated in Table 8.
@@ 2.2.7 GTC
The 10.4 m Gran Telescopio CANARIAS (GTC) telescope, at the Roque de los Muchachos Observatory, on the Island of La Palma (Spain) observed the afterglow of GRB 240825A. The observation started on August 26 at 02:04:10 UT, 10.190 hr after the burst onset. It was performed with the OSIRIS+ (Cepa et al. 2000) and consisted of s acquisition images in -band followed by s spectra using grism R1000B, covering the range between 3600 and 7800 Å, with a slit of 1′′ aligned with the parallactic angle, providing a resolving power of R . The mean epoch of the spectroscopic observation was 02:27:08 UT, 10.569 hr after the burst. The spectrum displays a clear continuum with absorption features due to Fe ii, Mg ii, Mg i, and Ca ii and emission due to [O ii] at a common redshift of 0.65960.0006. The spectrum is shown in the top panel of Figure 5.
Table 1 displays the equivalent width measurements of the prominent absorption features measured in the GTC spectrum. These values are used to calculate the Line Strength Parameter (LSP), which allows us to compare the strength of the absorption features of a GRB spectrum with those of a general GRB line-of-sight sample, independently of the spectral resolution or wavelength coverage. Following the prescription of de Ugarte Postigo et al. (2012), we obtain a value of LSP , which implies that the lines in this spectrum are slightly stronger than those of the typical GRB sight-line, corresponding to the 72 percentile of the samples of Fynbo et al. (2009) and de Ugarte Postigo et al. (2012) (the lines are stronger than those of 72% of the sample). The bottom panel of Figure 5 shows the Line Strength Diagram, which shows the comparison with the sample in more detail, comparing individual features. The line in black represents the average strength of the lines in the sample, and the dotted lines mark the standard deviation from that value. The red dots are the measurements from our spectrum. In this case, the Fe ii lines are almost equal to the sample, whereas the Mg ii and Mg i features are 1- stronger and the sample. The Ca ii are slightly stronger than those of the sample.
@@
Feii2374 | 33886.3 | (2.951.23)∗ |
Feii2383 | 33952.7 | 3.841.15 |
Feii2587 | 34293.6 | 2.340.75 |
Feii2600 | 34316.5 | 2.990.77 |
Mgii/Mgii2800 | 34647.4 | 11.170.69 |
Mgi2852 | 34737.0 | 3.960.47 |
Caii3935 | 36529.5 | 3.080.26 |
Caii3970 | 36586.5 | 2.200.27 |
@@ 2.2.8 VLT
We also carried out observations of the optical counterpart using the X-shooter instrument (Vernet et al. 2011), a multi-wavelength medium-resolution spectrograph installed at the European Southern Observatory (ESO) Very Large Telescope (VLT) UT3 (Melipal) under the program 110.24CF (PI: N. Tanvir). Our sequence started on 2024-08-26 at 02:18:03 (10.42 hours after the trigger) with photometric measurements of the optical counterpart in g′, r′, and z′ SDSS filters using the acquisition and guiding camera of the X-shooter (see Table 8). The spectroscopic observations were then obtained at a mid-time of 11.35 hours after the Swift trigger. Our spectra cover the wavelength range Å, and consisted of 4 exposures of 1200 s each. The full data set were reduced using the ESO pipeline following the strategy detailed in Goldoni et al. (2006) and Modigliani et al. (2010). Wavelengths were corrected to the vacuum-heliocentric system.
In the calibrated spectrum, a trace due to the afterglow emission is well detected over the entire wavelength range and several absorption features were identified due to Fe ii, Fe ii*, Mn ii, Mg ii, Mg i, Ca ii, Ca i and Na i. Using all these metal absorption lines, we measured as the redshift of the GRB (Martin-Carrillo et al. 2024). In addition, at , we also detect multiple bright emission lines produced by the GRB host galaxy as the [O II] and [O III] doublets, H, and H, together with [N II] and [S II] doublets (partially affected by sky lines). We note that the small shift between the redshift measured from the emission and absorption lines is consistent within the 1 errors and corresponds to a velocity shift of km s-1 towards the blue of the emission lines with respect to the absorption lines. The properties of the host galaxy will be investigated in more detail in a future work (Schneider et al., in prep.).
@@ 2.2.9 Asiago/REM/TNG/LBT
We obtained an optical Sloan- band observation with the 0.67/0.92 m robotic Schmidt telescope, located on Mount Ekar at the Asiago Observatory (Italy). The observation started on August 25, 2024 at 21:39:05 UT, which is 5.8 hours after the GRB trigger and lasted for 20 minutes. The data were reduced using the Snoopy pipeline141414Snoopy is a package for SN photometry using PSF fitting and/or template subtraction developed by E. Cappellaro. A package description can be found at http://sngroup.oapd.inaf.it/snoopy.html., which was used also to perform PSF photometry, calibrated against the Pan-STARRS catalogue. The astrometry was calibrated against field stars in the GAIA DR3 catalogue. The measured magnitudes are listed in Table 9.
@@
We obtained optical/NIR observations with the 0.6 m robotic Rapid Eye Mount telescope (REM, Zerbi et al. 2001; Covino et al. 2004), located at the European Southern Observatory (ESO) at La Silla (Chile). The observations started on August 26, 2024 at 00:28:43 UT, which is 8.6 hours after the GRB trigger and lasted for about 1 hour. Data reduction was performed automatically by the REM reduction pipeline: after bias subtraction, non-uniformities were corrected using a normalized flat-field frame processed with tools from the Swift Reduction Package (SRP)151515http://www.me.oa-brera.inaf.it/utenti/covino/usermanual.html. NIR data were sky-subtracted using the median of individual frames. Frame registration was performed using the Python-based software Astroalign (Beroiz et al. 2020), and astrometric solutions were derived against GAIA DR3 stars (Gaia Collaboration et al. 2023b). An uncatalogued source was detected, at a position consistent with the optical afterglow, in the stacked - and -band images. Aperture photometry was performed on the candidate with the SExtractor package (Version 2.28.1 Bertin & Arnouts 1996) and results were calibrated against Pan-STARRS DR2 (Flewelling 2018b) and 2MASS (Skrutskie et al. 2006) sources for the optical and NIR images, respectively.
@@
The field of GRB 240825A was also observed with the DOLoRes camera in spectroscopic mode mounted on the 3.6 m Telescopio Nazionale Galileo (TNG) at La Palma, Spain. Observations were carried out with the LR-B grism, covering the range 3500-8000 Å, with slit. The observations consisted of a spectrum carried out at a mean time of 7.4 hours after the burst. The optical afterglow is detected in the acquisition image, and aperture photometry was carried out with the SExtractor package and calibrated against the Pan-STARRS DR2 catalogue. The spectrum was reduced following standard procedure with the IRAF package (Version 2.16.1 Tody 1986), including bias subtraction, flat-field correction, wavelength calibration, and extraction. Despite the modest SNR we identified a single absorption feature consistent with Mg II at a redshift of (Figure 6), in agreement with the results reported by Martin-Carrillo et al. (2024). Further observations were performed with the Near Infrared Camera Spectrometer (NICS) mounted on the TNG telescope (Baffa et al. 2001) at a mid-time d. Data were reduced and stacked together with the jitter tool of the eclipse package, and the NIR afterglow was clearly detected. Aperture photometry was performed with SExtractor and calibrated against the 2MASS catalogue. We also obtained -band observations with the DOLoRes camera mounted on the TNG at a mid-time of about 13.4 days from the GRB trigger. The images were reduced and sky-subtracted with the jitter tool of the eclipse package. The source is visible at the position of the optical afterglow, aperture photometry was performed with the SExtractor package and calibration was against the Pan-STARRS DR2 catalogue. All measured magnitudes are listed in Table 9.
@@
On November 9, 2024, we also imaged the host in the -bands with the Large Binocular Cameras (LBCs; Giallongo et al. 2008) mounted on the Large Binocular Telescope (LBT) on Mt. Graham, Arizona, USA. The observations started at 03:58 UT, for a total of 3000 s in each band. LBT imaging data were reduced using the dedicated data reduction pipeline (Fontana et al. 2014). The astrometry was calibrated against field stars in the GAIA DR3 catalogue. We performed aperture photometry using DAOPHOT and APPHOT under PyRAF/IRAF and results were calibrated against the Pan-STARRS survey. All measured magnitudes are listed in Table 9.
@@ 2.2.10 MISTRAL
We finally carried out additional observations of the optical afterglow using the MISTRAL instrument mounted on the 193 cm telescope at Observatoire de Haute-Provence (OHP, France). The mid-time of these observations was 2024 August 25, 22:40 UT, corresponding to 6.85 hr after the GRB trigger. Three exposures of 5 min were obtained in the SDSS r′ band, leading to a total exposure time of 15 min. Data were reduced using standard bias- and flat-removal procedures, and the photometry was calibrated against nearby stars from the Pan-STARRS catalog. The magnitude we derived is listed in Table 8.
@@ 3.1 Afterglow model
The interaction between the relativistic outflow of the GRB and the circumburst medium can drive a pair of relativistic shocks: a forward shock (FS) and a reverse shock (RS). Electrons accelerated by these shocks produce multi-wavelength emission through synchrotron radiation and inverse Compton scattering. We will analyze the afterglow data in the NIR, optical, X-ray, and GeV bands given in Section 2.1.2 and Section 2.2, as well as the photometric data from Mephisto (Cheng et al. 2025).
@@
For afterglow modeling, numerous paradigms for numerical calculations have been established, generally starting from shock dynamical evolution (Huang et al. 1999; Nava et al. 2013; Zhang 2018), solving the electron continuity equation (Sari et al. 1998; Sari & Esin 2001; Fan & Piran 2006; Fan et al. 2008; Kumar et al. 2012), calculating the synchrotron radiation (Sari et al. 1998; Sari & Piran 1999) and synchrotron self-Compton (SSC; e.g. Fan et al. 2008; Nakar et al. 2009; Geng et al. 2018) scattering of electrons, and considering the impacts of various effects, including Klein-Nishina (KN; e.g. Fan et al. 2008; Nakar et al. 2009) effects, annihilation effects (e.g. Gould & Schréder 1967; Murase et al. 2011; Geng et al. 2018; Huang 2022), synchrotron self-absorption (SSA; e.g. Granot et al. 1999), as well as geometric and observational effects (Waxman 1997; Geng et al. 2018). Here, we use the Python-wrapped Fortran package ASGARD to perform the numerical calculations of the FS radiation under the various effects mentioned above (Ren et al. 2024). The calculation of the reverse shock (RS) contribution in the early afterglow follows the analytical method described in Yi et al. (2013), which includes only synchrotron emission. Given the large uncertainties and limited number of data points in the GeV energy band, we neglect the SSC contribution. This simplification may to some extent affect the estimation of the characteristic frequencies in the synchrotron emission. In this work, we assume that the jet has a “Top-Hat” structure. Therefore, under this model, the flux density at a certain time and frequency is:
where is the initial Lorentz factor, () is the fraction of the shock energy density converted into energy of relativistic electrons (magnetic field), is the half-opening angle in radians, is the isotropic kinetic energy, is the electron energy distribution index, is the medium number density. The additional subscripts and represent the FS and RS, respectively. Additionally, the host galaxy contribution is derived from the SED fitting results presented by Cheng et al. (2025). The model also accounts for several additional factors. For example, for the optical data, it adopts the Fitzpatrick (1999) extinction law to calculate both Galactic and host galaxy extinction. The Galactic extinction parameters are obtained from the IRSA Dust Extinction Service 161616https://irsa.ipac.caltech.edu/applications/DUST/, with = 3.1 and = 0.053. For the host galaxy extinction, the as the SMC is assumed (Kann et al. 2006; Schady et al. 2010), while is treated as a free parameter in the parameter inference.

### v2

@@ 2.2.5 LCOGT
The Las Cumbres Observatory Global Telescope Network (LCOGT; Brown et al. 2013) observed GRB 240825A using the 1 m telescope. Observations of the afterglow of GRB 240825A began at 01:41:29 UT on August 26, 2024, 9.81 hr after the GBM trigger. A total of 10 observing runs were conducted between August 26 and 31, resulting in a set of B, V, R, and I band images. The exposure time of each image was set to 300 s, based on the trend of brightness estimated from other observations’ measurements and the limiting magnitude of the LCOGT telescopes. The photometry calibration is carried out using the USNO-B1 catalogs (Monet et al. 2003). The source is clearly detected by stacking the images from the first nine observations. All photometric data are tabulated in Table 7.
@@ 2.2.6 NOT
We triggered the 2.56 m Nordic Optical Telescope (NOT) at the Roque de los Muchachos Observatory, La Palma, Spain, several times, utilizing the Alhambra Faint Object Spectrograph and Camera (ALFOSC). Photometric calibration was performed using the Pan-STARRS DR2 catalog (Flewelling 2018a). Photometric data are tabulated in Table 8.
@@ 2.2.7 GTC
The 10.4 m Gran Telescopio CANARIAS (GTC) telescope, at the Roque de los Muchachos Observatory, on the Island of La Palma (Spain) observed the afterglow of GRB 240825A. The observation started on August 26 at 02:04:10 UT, 10.190 hr after the burst onset. It was performed with OSIRIS+ (Cepa et al. 2000) and consisted of s acquisition images in -band followed by s spectra using grism R1000B, covering the range between 3600 and 7800 Å, with a slit of 1′′ aligned with the parallactic angle, providing a resolving power of R . The mean epoch of the spectroscopic observation was 02:27:08 UT, 10.569 hr after the burst. The spectrum displays a clear continuum with absorption features due to Fe II, Mg II, Mg I, and Ca II and emission due to [O II] at a common redshift of 0.65960.0006. The spectrum is shown in the top panel of Figure 5.
Table 1 displays the equivalent width measurements of the prominent absorption features measured in the GTC spectrum. These values are used to calculate the line strength parameter (LSP), which allows us to compare the strength of the absorption features of a GRB spectrum with those of a general GRB line-of-sight sample, independently of the spectral resolution or wavelength coverage. Following the prescription of de Ugarte Postigo et al. (2012), we obtain a value of LSP , which implies that the lines in this spectrum are slightly stronger than those of the typical GRB sight-line, corresponding to the 72 percentile of the samples of Fynbo et al. (2009) and de Ugarte Postigo et al. (2012) (the lines are stronger than those of 72% of the sample). The bottom panel of Figure 5 shows the Line Strength Diagram, which shows the comparison with the sample in more detail, comparing individual features. The line in black represents the average strength of the lines in the sample, and the dotted lines mark the standard deviation from that value. The red dots are the measurements from our spectrum. In this case, the Fe II lines are almost equal to the sample, whereas the Mg II and Mg I features are 1- stronger than the sample. The Ca II lines are slightly stronger than those of the sample.
@@
| FeII2374 | 3937.5 | (2.951.23)∗ |
| FeII2383 | 3952.7 | 3.841.15 |
| FeII2587 | 4293.6 | 2.340.75 |
| FeII2600 | 4316.5 | 2.990.77 |
| MgII2800 blend | 4647.4 | 11.170.69 |
| MgI2852 | 4737.0 | 3.960.47 |
| CaII3935 | 6529.5 | 3.080.26 |
| CaII3970 | 6586.5 | 2.200.27 |
@@ 2.2.8 VLT
We also carried out observations of the optical counterpart using the X-shooter instrument (Vernet et al. 2011), a multiwavelength medium-resolution spectrograph installed at the European Southern Observatory (ESO) Very Large Telescope (VLT) UT3 (Melipal) under the program 110.24CF (PI: N. Tanvir). Our sequence started on 2024 August 26 at 02:18:03 (10.42 hr after the trigger) with photometric measurements of the optical counterpart in g′, r′, and z′ SDSS filters using the acquisition and guiding camera of the X-shooter (see Table 8). The spectroscopic observations were then obtained at a mid-time of 11.35 hr after the Swift trigger. Our spectra cover the wavelength range Å, and consisted of four exposures, each 1200 s. The full data set was reduced using the ESO pipeline following the strategy detailed in Goldoni et al. (2006) and Modigliani et al. (2010). Wavelengths were corrected to the vacuum-heliocentric system.
In the calibrated spectrum, a trace due to the afterglow emission is well detected over the entire wavelength range and several absorption features were identified due to Fe II, Fe II*, Mn II, Mg II, Mg I, Ca II, Ca I and Na I. Using all these metal absorption lines, we measured as the redshift of the GRB (Martin-Carrillo et al. 2024). In addition, at , we also detect multiple bright emission lines produced by the GRB host galaxy as the [O II] and [O III] doublets, H, and H, together with [N II] and [S II] doublets (partially affected by sky lines). We note that the small shift between the redshift measured from the emission and absorption lines is consistent within the 1 errors and corresponds to a velocity shift of km s-1 towards the blue of the emission lines with respect to the absorption lines. The properties of the host galaxy will be investigated in more detail in a future work (Schneider et al., in preparation).
@@ 2.2.9 Asiago/REM/TNG/LBT
We obtained an optical Sloan- band observation with the 0.67/0.92 m robotic Schmidt telescope, located on Mount Ekar at the Asiago Observatory (Italy). The observation started on 2024 August 25 at 21:39:05 UT, which is 5.8 hr after the GRB trigger and lasted for 20 minutes. The data were reduced using the Snoopy pipeline141414Snoopy is a package for SN photometry using point-spread function (PSF) fitting and/or template subtraction developed by E. Cappellaro. A package description can be found at http://sngroup.oapd.inaf.it/snoopy.html., which was used also to perform PSF photometry, calibrated against the Pan-STARRS catalog. The astrometry was calibrated against field stars in the Gaia DR3 catalog. The measured magnitudes are listed in Table 9.
@@
We obtained optical/near-infrared (NIR) observations with the 0.6 m robotic Rapid Eye Mount telescope (REM, Zerbi et al. 2001; Covino et al. 2004), located at the ESO at La Silla (Chile). The observations started on 2024 August 26 at 00:28:43 UT, which is 8.6 hr after the GRB trigger and lasted for about 1 hour. Data reduction was performed automatically by the REM reduction pipeline: after bias subtraction, non-uniformities were corrected using a normalized flat-field frame processed with tools from the Swift Reduction Package (SRP)151515http://www.me.oa-brera.inaf.it/utenti/covino/usermanual.html. NIR data were sky-subtracted using the median of individual frames. Frame registration was performed using the Python-based software Astroalign (Beroiz et al. 2020), and astrometric solutions were derived against Gaia DR3 stars (Gaia Collaboration et al. 2023b). An uncataloged source was detected, at a position consistent with the optical afterglow, in the stacked - and -band images. Aperture photometry was performed on the candidate with the SExtractor package (Version 2.28.1 Bertin & Arnouts 1996) and results were calibrated against Pan-STARRS DR2 (Flewelling 2018b) and 2MASS (Skrutskie et al. 2006) sources for the optical and NIR images, respectively.
@@
The field of GRB 240825A was also observed with the DOLoRes camera in spectroscopic mode mounted on the 3.6 m Telescopio Nazionale Galileo (TNG) at La Palma, Spain. Observations were carried out with the LR-B grism, covering the range 3500-8000 Å, with slit. The observations consisted of a spectrum carried out at a mean time of 7.4 hr after the burst. The optical afterglow is detected in the acquisition image, and aperture photometry was carried out with the SExtractor package and calibrated against the Pan-STARRS DR2 catalog. The spectrum was reduced following standard procedure with the IRAF package (Version 2.16.1 Tody 1986), including bias subtraction, flat-field correction, wavelength calibration, and extraction. Despite the modest SNR we identified a single absorption feature consistent with Mg II at a redshift of (Figure 6), in agreement with the results reported by Martin-Carrillo et al. (2024). Further observations were performed with the Near Infrared Camera Spectrometer (NICS) mounted on the TNG telescope (Baffa et al. 2001) at a mid-time d. Data were reduced and stacked together with the jitter tool of the eclipse package, and the NIR afterglow was clearly detected. Aperture photometry was performed with SExtractor and calibrated against the 2MASS catalog. We also obtained -band observations with the DOLoRes camera mounted on the TNG at a mid-time of about 13.4 days from the GRB trigger. The images were reduced and sky-subtracted with the jitter tool of the eclipse package. The source was visible at the position of the optical afterglow, aperture photometry was performed with the SExtractor package and calibration was against the Pan-STARRS DR2 catalog. All measured magnitudes are listed in Table 9.
@@
On 2024 November 9, we also imaged the host in the -bands with the Large Binocular Cameras (LBCs; Giallongo et al. 2008) mounted on the Large Binocular Telescope (LBT) on Mt. Graham, Arizona, USA. The observations started at 03:58 UT, for a total of 3000 s in each band. LBT imaging data were reduced using the dedicated data reduction pipeline (Fontana et al. 2014). The astrometry was calibrated against field stars in the Gaia DR3 catalog. We performed aperture photometry using DAOPHOT and APPHOT under PyRAF/IRAF and results were calibrated against the Pan-STARRS survey. All measured magnitudes are listed in Table 9.
@@ 2.2.10 MISTRAL
We finally carried out additional observations of the optical afterglow using the MISTRAL instrument mounted on the 193 cm telescope at Observatoire de Haute-Provence (OHP, France). The mid-time of these observations was 2024 August 25, 22:40 UT, corresponding to 6.85 hr after the GRB trigger. Three exposures of 5 minutes were obtained in the SDSS r′ band, leading to a total exposure time of 15 minutes. Data were reduced using standard bias- and flat-removal procedures, and the photometry was calibrated against nearby stars from the Pan-STARRS catalog. The magnitude we derived is listed in Table 9.
@@ 3.1 Afterglow model
The interaction between the relativistic outflow of the GRB and the circumburst medium can drive a pair of relativistic shocks: an FS and an RS. Electrons accelerated by these shocks produce multiwavelength emission through synchrotron radiation and inverse Compton scattering. We will analyze the afterglow data in the NIR, optical, X-ray, and GeV bands given in Sections 2.1.2 and 2.2, as well as the photometric data from Mephisto (Cheng et al. 2025).
@@
For afterglow modeling, numerous paradigms for numerical calculations have been established, generally starting from shock dynamical evolution (Huang et al. 1999; Nava et al. 2013; Zhang 2018), solving the electron continuity equation (Sari et al. 1998; Sari & Esin 2001; Fan & Piran 2006; Fan et al. 2008; Kumar et al. 2012), calculating the synchrotron radiation (Sari et al. 1998; Sari & Piran 1999) and synchrotron self-Compton (SSC; e.g. Fan et al. 2008; Nakar et al. 2009; Geng et al. 2018) scattering of electrons, and considering the impacts of various effects, including Klein-Nishina (KN; e.g. Fan et al. 2008; Nakar et al. 2009) effects, annihilation effects (e.g. Gould & Schréder 1967; Murase et al. 2011; Geng et al. 2018; Huang 2022), synchrotron self-absorption (SSA; e.g. Granot et al. 1999), as well as geometric and observational effects (Waxman 1997; Geng et al. 2018). Here, we use the Python-wrapped Fortran package ASGARD to perform the numerical calculations of the FS radiation under the various effects mentioned above (Ren et al. 2024). The calculation of the RS contribution in the early afterglow follows the analytical method described in Yi et al. (2013), which includes only synchrotron emission. Given the large uncertainties and limited number of data points in the GeV energy band, we neglect the SSC contribution. This simplification may to some extent affect the estimation of the characteristic frequencies in the synchrotron emission. In this work, we assume that the jet has a “top-hat” structure. Therefore, under this model, the flux density at a certain time and frequency is:
where is the initial Lorentz factor, () is the fraction of the shock energy density converted into energy of relativistic electrons (magnetic field), is the half-opening angle in radians, is the isotropic kinetic energy, is the electron energy distribution index, and is the medium number density. The additional subscripts and represent the FS and RS, respectively. Additionally, the host galaxy contribution is derived from the SED fitting results presented by Cheng et al. (2025). The model also accounts for several additional factors. For example, for the optical data, it adopts the Fitzpatrick (1999) extinction law to calculate both Galactic and host galaxy extinction. The Galactic extinction parameters are obtained from the IRSA Dust Extinction Service 161616https://irsa.ipac.caltech.edu/applications/DUST/, with = 3.1 and = 0.053. For the host galaxy extinction, the of the SMC is assumed (Kann et al. 2006; Schady et al. 2010), while is treated as a free parameter in the parameter inference.